OpenStudy (crashonce):

if (5a-3b)/(a+b) = 3 then (a+b):(a-b) = ?

OpenStudy (crashonce):

@perl

OpenStudy (perl):

try to simplify (5a-3b)/(a+b) = 3

OpenStudy (crashonce):

a=0?

OpenStudy (anonymous):

What do you get when you multiply both sides by \(a+b\)?

OpenStudy (crashonce):

5a-3b = 3a+3b 2a=6b a=3b

OpenStudy (crashonce):

lol woops

OpenStudy (crashonce):

right its 2:1 sorry

OpenStudy (anonymous):

Wonderful